# Multi-weather Cross-view Geo-localization Using Denoising Diffusion Models

### v1

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: xxxxxxxxxisbn: xxxxxxxxxccs: Computing methodologies Image representationsccs: Information systems Top-k retrieval in databases
@@ 1. Introduction
Cross-view geo-localization(4, ) aims to determine an unknown location by matching drone-view images with the correct geo-tagged satellite-view images from a large gallery, based on geographic features in the images, as shown in Figure 1. This task is crucial for accurate navigation and safe planning(6, ; 7, ; 8, ) in GNSS-denied autonomous drone flights. Recent advances in vision transformer have led to significant breakthroughs in various cross-view geo-localization tasks, such as drone localization(1, ; 5, ) (matching drone-view query images with geo-tagged satellite-view images) and drone navigation(2, ; 3, ) (using satellite-view query images to guide drones to a target area). However, varying weather conditions, including fog, rain, snow, wind, light, dark, and combinations of multiple weather types, reduce visibility, corrupt the information captured by an image, significantly complicate image geographic representation, and lead to a sharp decline in task performance. The major challenge lies in adaptively achieving unbiased image geographic representation under diverse weather conditions.
@@
A clean image without any weather degradation is desired in cross-view geo-localization. Early methods for weather removal using empirical observations (9, ), Convolutional Neural Networks (CNNs) based and transformer-based for deraining(10, ), dehazing(11, ), and desnowing(12, ). Most of these methods achieve excellent performance, but these are not generic solutions for all adverse weather removal problems as the networks have to be trained separately for each weather(13, ). The All-in-One Network(14, ) proposes a framework with separate encoders for each weather but a generic decoder and neural architecture search across weather-specific optimized encoders. The Transweather(15, ) using vision transformer construction has a single encoder and a decoder and learns weather-type queries to solve all adverse weather removal efficiently. Wetherdiff(16, ) using diffusion models enables size-agnostic image restoration by using a guided denoising process. To our interest, these three studies focus on the inability of specific weather combinations to adapt to new weather types. Recently, MuSe-Net(17, ) employs a two-branch neural network containing one multiple-environment style extraction network and one self-adaptive feature extraction network to dynamically adjust the domain shift caused by environmental changes. However, this method does not perform well in some real-world high-intensity rains with a splattering effect. In summary, multi-weather cross-view geo-localization in unseen unpredictable real weather conditions is a problem that needs to be solved urgently.
To overcome these obstacles, this paper presents MCGF, a Multi-weather Cross-view Geo-localization Framework designed to dynamically adapt to unseen weather conditions, which establishes a joint optimization between image restoration and geo-localization using denoising diffusion models. Diffusion models increasingly serve discriminative tasks such as classification and image segmentation. Inspired by its powerful modeling capability and stable training process, we utilize the diffusion model to learn the denoising process from noisy images to clean images, facilitating robust matching in the presence of multi-weather. In image restoration, MCGF includes a shared encoder and a lightweight restoration module that prompts the backbone to provide more beneficial information to eliminate the influence of weather-specific information. In geo-localization, MCGF uses EVA-02(18, ) as a backbone for feature extraction and uses cross-entropy loss for training and cosine distance for testing. EVA-02 is a ViT(19, ) model obtained using a series of stable optimization methods, which allows MCGF to extract more favorable information from drone and satellite images while using fewer parameters.
@@
Figure 2. The overview of MCGF, which establishes a joint optimization between image restoration and geo-localization.
@@ 2.2. Shared Encoder
To enhance feature representation and improve subsequent image restoration and geo-localization, we utilize the widely adopted state-of-the-art transformer-based model, Swin Transformer(21, ), as the shared encoder in our unified framework. The Swin Transformer is a hierarchical transformer that employs shifted windows, which restricts attention computation to non-overlapping local windows, making it adaptable for modeling at various scales. To balance computational overhead and inference speed, we select the tiny version of Swin Transformer as the default backbone.
@@ 3. Experiment
Implement details. We employed the EVA-02 model, which is based on the Vision Transformer, as the backbone for diffusive matching module. This model has been trained and fine-tuned on many large vision datasets. In our experiments, we resized each input image to a fixed size of 448 × 448 pixels. During training, we used SGD as the optimizer with a momentum of 0.9 and weight decay of 5 × $10^{(}-4)$, with a mini-batch size of 16. The initial learning rate was set to 0.01 for the backbone layer and 0.1 for the classification layer. Our model was built using Pytorch.
@@ 3.1. Geo-localization results
We train MCGF with outstanding algorithms (including LPN(22, ), MBEG(23, ), and Muse-Net(17, )) on the University-160k-WX train set until convergence and obtain optimal results. We test all trained models on the official unified test set provided by the competition organizer. All test results can be displayed and downloaded on the competition result submission platform. Table 1 shows that MCGF is significantly better than existing methods in all evaluation metrics. Especially compared with the latest research Muse-Net, MCGF can achieve a 67.75% performance improvement in the Recall@1 indicator. MCGF shows considerable potential for geo-localization as a general framework.
@@ 5. Acknowledgments
LPN(22, ) 7.98 10.25 11.21 8.49
MBEG(23, ) 26.17 32.84 35.32 29.32
Muse-Net(17, ) 50.48 63.19 67.34 53.27

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2nd Workshop on UAVs in Multimedia: Capturing the World from a New Perspective; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 2nd Workshop on UAVs in Multimedia: Capturing the World from a New Perspective (UAVM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3689095.3689103isbn: 979-8-4007-1206-7/24/10ccs: Computing methodologies Image representationsccs: Information systems Top-k retrieval in databases
@@ 1. Introduction
Cross-view geo-localization(Shi et al., 2022a) aims to determine an unknown location by matching drone-view images with the correct geo-tagged satellite-view images from a large gallery, based on geographic features in the images, as shown in Figure 1. This task is crucial for accurate navigation and safe planning(Wei et al., 2024; Lu et al., 2024; Zhao et al., 2023) in GNSS-denied autonomous drone flights. Recent advances in vision transformer have led to significant breakthroughs in various cross-view geo-localization tasks, such as drone localization(Song et al., 2024; Shi et al., 2022b) (matching drone-view query images with geo-tagged satellite-view images) and drone navigation(Shi et al., 2023; Shi and Li, 2022) (using satellite-view query images to guide drones to a target area). However, varying weather conditions, including fog, rain, snow, wind, light, dark, and combinations of multiple weather types, reduce visibility, corrupt the information captured by an image, significantly complicate image geographic representation, and lead to a sharp decline in task performance. The major challenge lies in adaptively achieving unbiased image geographic representation under diverse weather conditions.
@@
A clean image without any weather degradation is desired in cross-view geo-localization. Early methods for weather removal using empirical observations (He et al., 2010a), Convolutional Neural Networks (CNNs) based and transformer-based for deraining(He et al., 2010b), dehazing(Zhang et al., 2021b), and desnowing(Zhang et al., 2021a). Most of these methods achieve excellent performance, but these are not generic solutions for all adverse weather removal problems as the networks have to be trained separately for each weather(Zhao et al., 2021). The All-in-One Network(Li et al., 2020) proposes a framework with separate encoders for each weather but a generic decoder and neural architecture search across weather-specific optimized encoders. The Transweather(Valanarasu et al., 2022) using vision transformer construction has a single encoder and a decoder and learns weather-type queries to solve all adverse weather removal efficiently. Wetherdiff(Özdenizci and Legenstein, 2023) using diffusion models enables size-agnostic image restoration by using a guided denoising process. To our interest, these three studies focus on the inability of specific weather combinations to adapt to new weather types. Recently, MuSe-Net(Wang et al., 2024) employs a two-branch neural network containing one multiple-environment style extraction network and one self-adaptive feature extraction network to dynamically adjust the domain shift caused by environmental changes. However, this method does not perform well in some real-world high-intensity rains with a splattering effect. In summary, multi-weather cross-view geo-localization in unseen unpredictable real weather conditions is a problem that needs to be solved urgently.
To overcome these obstacles, this paper presents MCGF, a Multi-weather Cross-view Geo-localization Framework designed to dynamically adapt to unseen weather conditions, which establishes a joint optimization between image restoration and geo-localization using denoising diffusion models. Diffusion models increasingly serve discriminative tasks such as classification and image segmentation. Inspired by its powerful modeling capability and stable training process, we utilize the diffusion model to learn the denoising process from noisy images to clean images, facilitating robust matching in the presence of multi-weather. In image restoration, MCGF includes a shared encoder and a lightweight restoration module that prompts the backbone to provide more beneficial information to eliminate the influence of weather-specific information. In geo-localization, MCGF uses EVA-02(Fang et al., 2023) as a backbone for feature extraction and uses cross-entropy loss for training and cosine distance for testing. EVA-02 is a ViT(Dosovitskiy et al., 2020) model obtained using a series of stable optimization methods, which allows MCGF to extract more favorable information from drone and satellite images while using fewer parameters.
@@
Figure 2. The overview structure of MCGF, which establishes a joint optimization between image restoration and geo-localization.
@@ 2.2. Shared Encoder
To enhance feature representation and improve subsequent image restoration and geo-localization, we utilize the widely adopted state-of-the-art transformer-based model, Swin Transformer(Liu et al., 2021), as the shared encoder in our unified framework. The Swin Transformer is a hierarchical transformer that employs shifted windows, which restricts attention computation to non-overlapping local windows, making it adaptable for modeling at various scales. To balance computational overhead and inference speed, we select the tiny version of Swin Transformer as the default backbone.
@@ 3. Experiment
Implement details. We employed the EVA-02 model, which is based on the Vision Transformer, as the backbone for diffusive matching module. This model has been trained and fine-tuned on many large vision datasets. In our experiments, we resized each input image to a fixed size of 448 × 448 pixels. During training, we used SGD as the optimizer with a momentum of 0.9 and weight decay of 5 × $10^{-4}$, with a mini-batch size of 16. The initial learning rate was set to 0.01 for the backbone layer and 0.1 for the classification layer. Our model was built using Pytorch.
@@ 3.1. Geo-localization results
We train MCGF with outstanding algorithms (including LPN(Wang et al., 2021), MBEG(Zhu et al., 2023), and Muse-Net(Wang et al., 2024)) on the University-160k-WX train set until convergence and obtain optimal results. We test all trained models on the official unified test set provided by the competition organizer. All test results can be displayed and downloaded on the competition result submission platform. Table 1 shows that MCGF is significantly better than existing methods in all evaluation metrics. Especially compared with the latest research Muse-Net, MCGF can achieve a 67.75% performance improvement in the Recall@1 indicator. MCGF shows considerable potential for geo-localization as a general framework.
@@ 5. Acknowledgments
LPN(Wang et al., 2021) 7.98 10.25 11.21 8.49
MBEG(Zhu et al., 2023) 26.17 32.84 35.32 29.32
Muse-Net(Wang et al., 2024) 50.48 63.19 67.34 53.27